How many sides does a polygon have if its angle sum is 1080 degrees?
Answer by nyc_function(2741)

The interior angles of a polygon add up to 180(n-2).
We solve:

180(n-2) = 1080

180n - 360 = 1080

180n = 1080 + 360

180n = 1440

n = 1440/180

n = 8

The answer is 8 sides.